OpenStudy (anonymous):

i need help the first difference of a sequence are a constante 4 and the second term 8 what are the fisrt 5 terms of the sequence??

OpenStudy (anonymous):

@susannn .. see d= 4 and a=8 .. according to your question... Now, each term can be found out using the formula,,,,, t (n) = a + (n-1) d put, n = 1,2,3,4,5... and find all the terms... Good luck..

OpenStudy (anonymous):

i cant get it right !

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

tell me the answer.. and I will tell you if it is right .. let me do the first one for you... First term , for n=1, t(1) = a + (1-1) *d = a +0 *d = a = 8 Second term, for n = 2, t(2)= a + (2-1) * d = a + d = 8 +4 = 12 Now you can do the rest...

OpenStudy (anonymous):

thankzzz alot (Y)